# Building a Scalable, Effective, and Steerable Search and Ranking Platform

### v1

@@ 7. Model Productionization
For scoring items in the candidate generation phase, we utilize Elasticsearch’s vector search function, which employs an efficient approximate k-NN search. We also incorporate real-time customer action ingestion, which processes the resulting sequences to compute customer embeddings and feed the ranking layer. A caching layer positioned between these two processes stores the customer sequences, which can be uniformly fed to both models thanks to the uniformity of architecture. This ensures high throughput and maintains low latency during inference.
@@ 8. EXPERIMENTS
In this section, we present the results of extensive online (A/B testing) and offline experimentation and ablation studies using internal datasets. We compare the newly introduced ranking system against the existing baseline models.
@@ 8.1.1. Dataset
The offline dataset consists of a sample of item interactions from the last 60 days aggregated by customer id. The item interactions consist of product clicks, add-to-wishlist, add-to-cart, and checkout events attributed to the browse and search premise by using the ”last-touch” attribution model. The interactions are joined with the corresponding item ids, their timestamps, and interaction type and sorted by timestamp. To form a single data sample, the resulting sequences are combined with contextual data, specifically, market, device type, browsing category, and search query (if present). The training dataset consisted of 71M unique customers across 25 markets. We do not perform any preprocessing such as deduplication or outlier removal on the obtained customer sequences besides truncating to the last 100 actions. The average sequence length is 24 actions. The evaluation dataset contains 300K customers (Figure 5 in the Appendix provides histograms per action type). We applied hard temporal split to create the training and test datasets to ensure no data leakage.
@@ 8.1.2. Metrics and evaluation protocol
The main metrics that our models are evaluated on are the following:
  • Recall@k: proportion of all relevant items within top k items (as defined in (Tamm et al., 2021)).
  • NDCG@k: measures the effectiveness of a ranking, taking into account the position of relevant items in the ranked list of top-k items, and attributed items are considered as relevant (as defined in (Tamm et al., 2021)).
  • Diversity: we use the maximum run of consecutive items from the same brand as a proxy to brand diversity. A high value suggests that a ranking can be dominated by items coming from only a few brands. User acceptance testing showed this leads to undesirable customer experience.
@@ 8.1.3. Candidate Generation
  • GBT is a candidate generation model based on Gradient Boosting Trees, which has proven to offer competitive performance compared to neural-based models (Qin et al., 2021). It ranks items based on their static attributes (season, material, type, etc.) and their dynamic historical engagement rates (e.g. add-to-cart rate for the last 5 min). The LambdaRank objective (Burges et al., 2006) is used during training to up-rank interactions based on their graded relevance (”purchase” has the highest while ”click” the lowest relevance). The model’s scores are computed in streaming fashion making it highly reactive to trends in customer behavior. This model was our previous production model.
  • RCG is our new candidate generation model introduced in section 4. We test a few versions of this model: RCGntr which uses pre-trained visual embeddings of items and includes a bias term that captures item popularity; RCGtr which employs trainable item embeddings, initialized from the pre-trained visual item embeddings. These two models utilize only global contextual information, such as country and device type. Additionally, RCGtr+ctx denotes a variant with additional local contextual input — the current user’s browsing category and a binary flag whether they browse or search. All the model versions consist of 2 encoder layers with 4 heads, gelu activation, and a max. sequence length of 100. The model was trained for 20 epochs by using the Adam optimizer, with a learning rate set to 0.001. We do not consider hard negatives and rather employ log uniform candidate sampling.
Table 1. Ablation study of the ranking candidate generation model (RCG) on browse traffic relative to the baseline candidate generator GBT. Recall@500, which is the primary offline metric, is shown per customer segment while NDCG is shown for all customers.
@@
Table 1 presents an offline ablation study comparing our RCG model with the baseline GBT used as a candidate generator on the browse use case (Table 6 in the Appendix summarizes offline evaluation on the search use case). Across all evaluated customer segments, our newly proposed candidate generator significantly outperforms the existing one.
@@
Incorporating local contextual inputs further boosts the performance of the RCGtr model. Specifically, by including item category and search query presence, the RCGtr+ctx model’s Recall@500 is enhanced by an additional $10\%$, and its NDCG by $14\%$. In Section 8.2 we will show that these gains are also reflected in our online experiments.
@@ 8.1.4. Ranking Layer
Table 2. Relative performance change compared to WDL-ATT, on relevance, freshness, and diversity-oriented metrics.
@@
  • RL is our new ranking model introduced in section 5. $d_{\textrm{model}}$ and the model output is set to 128, the model consists of 2 encoder layers with 8 heads, relu activation, and a max. sequence length of 80. The model is trained for 2 epochs by using the Adam optimizer, with a learning rate of $0.001$. The ratio of negatives vs. positives is set to 4. Since sampling negatives only from items that were in the view-port lead to degradation of performance, we sampled from all non-interacted items in the page.
@@
It should be noted that all compared algorithms use the same near real-time serving infrastructure. Some of the other algorithms mentioned, such as TransAct (Xia et al., 2023), while potentially competitive, were not applicable to our use case or latency constraints. For simplicity, we used a single ranking model in the ranking layer. In these experiments, we focus on the NDCG metric for ”high-value actions” or HVAs, which, in our context, are add-to-wishlist and add-to-cart actions. This metric acts as a proxy for our success KPI, defined by customer engagement wrt. HVAs, as detailed in the subsequent section.
@@
Table 3 presents an ablation study that explains the contribution of each component and input type to the model’s performance. The removal of any of these elements significantly detracts from overall model efficacy. Particularly noticeable is the performance decline when contextual inputs are not integrated early in the encoder. This finding suggests that our model more effectively leverages contextual information alongside rich item representations in customer action sequences compared to other algorithms such as BST. We opted not to benchmark against the DCN method cited in (Xia et al., 2023), as self-attention offers a superior mechanism for capturing feature interactions within sequential data compared to feature crosses. Additionally, the ablation study demonstrates that omitting heterogeneous inputs markedly diminishes the model’s performance. Specifically, excluding contextual inputs results in more than a 10% decrease in NDCG@6, while complete removal of item metadata leads to a drastic 26% reduction in NDCG@6.
Table 3. Ablation study: relative change of the NDCG metric at $k=\{6,84\}$, measured after removing various components and/or inputs in the RL model architecture.
@@ 3 RCGtr + RL + PL - - - -
Table 5. Summary of A/B tests performed on the Search use case. Relative percent uplifts are reported. The best-performing variant in each test is highlighted in bold. Non-significant results are marked with *.
@@ 2 RCGtr + RL + PL - -
RCGtr+ctx + RL + PL 0.70% 0.17%
8.2. Online Experiments
We have conducted several online A/B tests on real-world ranking use cases at Zalando. These tests were carried out systematically, replacing one component at a time to evaluate its impact. All tests allocated equal traffic splits among variants over a few weeks, as necessary, to achieve the minimum detectable effect for the success KPI with a $p$-value$<0.05$. Each model was retrained and deployed daily. Beyond customer engagement, our evaluation of online performance encompasses a variety of exploratory metrics, including financial metrics, the capacity to promote new items (novelty), and ranking diversity (as defined in subsubsection 8.1.2 as well as in the number of distinct brands on the first page). Table 4 summarizes the results from a series of A/B tests performed on the browse use case, where a customer is browsing the category tree as described in Figure 1.
In the first A/B test in Table 4 (row 1), we compared the new candidate generation model (RCG) against the previous ranking system end-to-end (GBT + WDL-ATT). We examined two variants: firstly, the RCG model with trainable item embeddings (RCGtr) and, secondly, the RCG model with non-trainable item embeddings (RCGntr) paired with WDL-ATT as the ranker. The former variant demonstrated a significant increase in engagement. However, the new candidate generation model had no significant impact on financial KPIs, promoted fewer new items, and decreased diversity.
@@ 8.2. Online Experiments
Table 5 summarizes the results of a series A/B tests performed on the search use case, i.e., when the customer is using full-text search to find their desired item. The conclusions are similar and for brevity, we omit the details.
@@ 9. CONCLUSION
In this paper, we have introduced a flexible, scalable, steerable, and real-time ranking platform that has been proven to enhance customer experience by delivering more relevant and personalized content. This approach has, in turn, led to improvements in various customer-centric and business metrics. We have described the architecture of our ranking platform, adhering to a set of design principles and utilizing state-of-the-art models. We have also offered insights into their performance, highlighting the considerable advantages of integrating heterogeneous signals and inputs that encompass the entire customer journey, as well as the effectiveness of fine-tuning input embeddings to boost model performance.

### v2

@@ 7. Model Productionization
For scoring items in the candidate generation phase, we utilize ElasticSearch’s vector search function, which employs an efficient approximate k-NN search. An action ingestion is employed by using a near real-time feature store with a caching layer used to temporarily store the customer data ready to be fed to ranking models from different layers thanks to their similar model architecture. The product metadata maps are stored within the model; therefore, only IDs are transferred over the network. This improves the throughput and maintains low latency during inference.
@@ 8. EXPERIMENTS
In this section, we present the results of extensive online and offline experiments and ablation studies using internal datasets. We compare the newly introduced ranking system against the existing baseline models for browse, search and item recommendation use cases.
@@ 8.1.1. Dataset
The offline dataset consists of a sample of historical item interactions aggregated by customer id. The item interactions consist of product clicks, add-to-wishlist, add-to-cart, and checkout events attributed to the browse and search premise by using the ”last-touch” attribution model. These actions are used as implicit feedback for training and evaluation. The interactions are joined with the corresponding item IDs, their timestamps, and interaction type and sorted by timestamp. To form a single data sample, the resulting sequences are combined with contextual data, specifically, market, device type, browsing category, and search query (if present). The training dataset consisted of 71M unique customers across 25 markets. We do not perform any preprocessing such as de-duplication or outlier removal on the obtained customer sequences besides truncating to the last 100 actions. The average sequence length is 24 actions. The evaluation dataset contains 300K customers (Figure 5 in the Appendix provides histograms per action type). We applied hard temporal split to create the training and test datasets to ensure no data leakage.
@@ 8.1.2. Metrics and evaluation protocol
The ranking models have been evaluated on following metrics:
  • Recall@k: defined as the proportion of all relevant items within top k items (as defined in (Tamm et al., 2021));
  • NDCG@k: measures the effectiveness of a ranking by taking into consideration the position and the relevance label of each item in the ranked list of top-k items (as defined in (Tamm et al., 2021));
  • Diversity: we use the maximum run of consecutive items from the same brand as a proxy to brand diversity. A high value suggests that a ranking can be dominated by items coming from only a few brands. User acceptance testing showed this leads to undesirable customer experience;q
@@ 8.1.3. Candidate Generation
  • GBT is a candidate generation model based on Gradient Boosting Trees, which has proven to offer competitive performance compared to neural-based models (Qin et al., 2021). It ranks items based on their static metadata (season, material, category type, etc.) and their dynamic historical engagement rates (add-to-cart, add-to-wishlist and click rates for the last 5 min). The LambdaRank objective (Burges et al., 2006) is used during training to up-rank interactions based on their graded relevance (”purchase” has the highest while ”click” the lowest relevance). The model’s scores are computed in streaming fashion making it highly reactive to trends in customer behavior. This model was our previous production model.
  • RCG is our candidate generation model introduced in section 4. We test a few variants of this model: RCGntr which uses pre-trained visual embeddings of items (Bracher et al., 2016) and includes a bias term that captures item popularity; RCGtr which employs trainable item embeddings, initialized from the pre-trained visual item embeddings. These two models utilize only global contextual information, such as country and device type. Additionally, RCGtr+ctx denotes a variant with additional local contextual input — the current user’s browsing category and a binary flag whether they browse or search. All the model versions consist of 2 encoder layers with 4 heads, gelu activation, and a max. sequence length of 100. The model was trained for 20 epochs by using the Adam optimizer, with a learning rate set to 0.001. We do not consider hard negatives and rather employ log uniform candidate samplin as described in section 4.
Table 1. Ablation study of the ranking candidate generation model (RCG) on browse traffic relative to the baseline candidate generator GBT. Recall@500, which is the main offline metric for candidate generation, is shown per customer segment while NDCG is shown for all customers.
@@
Table 2. Performance summary of the ranking candidate generation (RCG) models on the search use case relative to the baseline candidate generation GBT model.
Model Recall@500 NDCG@500
RCGtr +2.7% +16.8%
RCGtr+ctx +4.6% +22.0%
Table 1 presents an offline ablation study comparing our RCG model with the baseline GBT used as a candidate generator on the browse use case while Table 2 summarizes the offline evaluation results of the RCG model on the search use case only. Similar improvements seen on the browse traffic apply to the search traffic. In summary, across all evaluated customer segments, our newly proposed candidate generator significantly outperforms the existing one.
@@
Incorporating local contextual inputs further boosts the performance of the RCGtr model. Specifically, by including item category and search query presence, the RCGtr+ctx model’s Recall@500 is enhanced by an additional $10\%$, and its NDCG by $14\%$. In subsection 8.2 we will show that these gains are also reflected in our online experiments.
@@ 8.1.4. Ranking Layer
Table 3. Relative performance change compared to WDL-ATT, on relevance, freshness, diversity and novelty.
@@
  • RL is our ranking model introduced in section 5. The model consists of 2 encoder layers with 8 heads, relu activation, and a maximum sequence length of 80. $d_{\textrm{model}}$ and the model output size is set to 128. The model is trained for 2 epochs by using the Adam optimizer, with a learning rate of $0.001$. The ratio of negative vs. positive samples is set to 4. Since sampling negatives only from items that were in the view-port lead to degradation of performance, we sampled from all non-interacted items on the page.
@@
It should be noted that all compared algorithms use the same near real-time serving infrastructure. Some of the algorithms mentioned in related work, such as TransAct (Xia et al., 2023), while potentially competitive, were not applicable to our use case due to latency constraints. In these experiments, we focus on the NDCG metric for ”high-value actions” or HVAs, which, in our context, are add-to-wishlist and add-to-cart actions. This metric acts as a proxy for our success KPI, defined by customer engagement wrt. HVAs which is described in more detail in subsection 8.2.
@@
Table 4 shows an ablation study that describes the contribution of each input type to the model’s performance. The removal of any of the inputs significantly affects the overall model accuracy. Particularly noticeable is the performance decline when contextual inputs are not integrated early in the encoder. This finding suggests that our model leverages contextual information more effectively alongside rich item representations in customer action sequences compared to algorithms such as BST that fail to capture complex interactions between contextual, action and item metadata. Additionally, the ablation study shows that omitting heterogeneous inputs substantially diminishes the model’s performance. Specifically, excluding contextual inputs results in more than a 10% decrease in NDCG@6, while complete removal of item metadata leads to a drastic 26% reduction in NDCG@6.
Table 4. Ablation study: relative change of the NDCG metric at $k=\{6,84\}$, measured after removing various inputs in the RL model.
@@ 3 RCGtr + RL + PL - - - -
Table 6. Summary of A/B tests performed on the search use case. The relative percent uplifts are reported with engagement being the success KPI. The best-performing variant in each test is highlighted in bold. Non-significant results are marked with *.
@@ 2 RCGtr + RL + PL - -
RCGtr+ctx + RL + PL +0.70% +0.17%
Table 7. Summary of A/B tests performed on recommendation use cases. The relative percent uplifts against the baseline are reported with engagement being the success KPI. The best-performing variant in each test is highlighted in bold.
Reco use case Variant Engagement
PDP (similar items) CF (baseline) -
CF + RL +0.78% HVA
Home (product campaign) Curated (baseline) -
RL +21.2% CTR
We have conducted several online A/B tests on real-world ranking and recommendation use cases at Zalando. Whenever possible, these tests were carried out systematically, replacing one component at a time to evaluate its impact. All tests allocated equal user splits among variants (a given users always remains in the initially allocated variant) over a few weeks, as necessary, to achieve the minimum detectable effect for the success KPI with a $p$-value$<0.05$. Each model was retrained and deployed daily. Beyond customer engagement, our evaluation of online performance sometimes includes exploratory metrics, including financial metrics, the capacity to promote new items (novelty), and ranking diversity (as defined in subsubsection 8.1.2).
Table 5 summarizes the results from a series of A/B tests performed on the browse use case, where a customer is browsing the category tree as depicted in Figure 1. In the first A/B test in Table 5 (row 1), we compared the new candidate generation model (RCG) against the previous ranking system end-to-end (GBT + WDL-ATT). We evaluated two variants: firstly, the RCG model with trainable item embeddings (RCGtr) and, secondly, the RCG model with non-trainable item embeddings (RCGntr) paired with WDL-ATT as the next ranker in the funnel. The former variant demonstrated a significant increase in engagement. However, the new candidate generation model had no significant impact on financial KPIs, promoted fewer new items, and decreased diversity.
@@
Table 6 summarizes the results of a series A/B tests performed on the search use case, i.e., when the customer is using full-text search to find their desired item(s). The conclusions are similar and for brevity we omit the details.
Table 7 summarizes the outcome of two A/B tests where RL was used to re-rank existing recommendation baselines. In the PDP (product detail page) use case, RL was used to re-rank the top-200 most similar items produced by k-NN item-to-item collaborative filtering (CF) out of which the top-15 are shown in the ”similar item” carousel. In the Home use case, RL was used to rerank a curated set of products from an active campaign and show these items in this order to the customer. As a conclusion, RL can be re-used to personalize other customer touch points.
8.3. Serving Latency
Processing of rich contextual and sequence information in the models can increase their latency, making them unsuitable in real-world applications. While we have seen an increase in the overall latency, the extra latency in our ranking comes from: i) introduction of the candidate generator endpoint, adding 10ms p99 latency on average for customer embeddings, and ii) a k-NN search in the vector database, adding 30ms p99 latency on average. Our new ranking model has similar inference latency to the legacy WDL-ATT model which is 10ms p99 on average. The total added latency was approximately 40ms, resulting in 200ms end-to-end latency, well below our 500ms SLO. All endpoints were deployed on standard CPU instances.
@@ 9. CONCLUSION
In this paper, we have introduced a flexible, scalable, steerable, and real-time ranking platform that has been proven to enhance customer experience by delivering more relevant and personalized content for various use cases including browse, search and item recommendation. The approach has led to improvements in both customer-centric and business metrics. We have described the architecture of our ranking platform, adhering to a set of design principles and utilizing state-of-the-art models. We have also provided insights into their performance, highlighting the considerable advantages of integrating heterogeneous signals and inputs that encompass the entire customer journey, as well as the effectiveness of fine-tuning input embeddings to boost model performance.